# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

### v3

@@ 1.3. Related works
Unlike many mixing time works which focus on the discrete-time finite state space Markov chains (Jerrum and Sinclair, 1989; Lovasz and Simonovits, 1990; Berestycki, 2020; Levin and Peres, 2017; Wilson, 2004), analyzing the mixing behavior of queuing systems is a non-trivial endeavor as it involves handling an infinite, countable state space. Indeed, while there have been prior works on the mixing time of $M/M/n$ queues (Robert, 2003; Gamarnik and Goldberg, 2013; Zeifman, 1991; van Leeuwaarden and Knessl, 2011) and other queuing systems (Lorek and Szekli, 2015; Mao and Xia, 2015; Gurvich, 2014; Braverman, 2020, 2023; Luczak and McDiarmid, 2006; Rutten and Mukherjee, 2023), each of them can either only obtain tight finite-time bounds that match the asymptotic behavior only for a specific regime or setting, or can only show exponential ergodicity without an explicit rate (Gurvich, 2014; Braverman, 2020, 2023), geometric convergence to a ball (Luczak and McDiarmid, 2006) or subgeometric convergence (Rutten and Mukherjee, 2023). In particular, we shall provide a recap of mixing results for each of these regimes and settings.
@@
$M/M/1$ queue: The paper (Morse, 1955) is the first to characterize the transient behavior of an $M/M/1$ queue. Following this work, (Van Doorn, 1985) uses the spectral representation of birth-and-death processes to obtain the rate of convergence. On the other hand, (Robert, 2003) uses a coupling argument to establish the distance to stationarity for the $M/M/1$ queue where the author upper bounds the distance to stationarity with the probability that the two processes will coalesce within time $T$, which can be further upper bounded by the probability of hitting 00 within time $T$. Unfortunately, both of these techniques are difficult to generalize to the $M/M/n$ queue given its more complicated dynamics.
@@
$M/M/\infty$ queue: The explicit characterization of the transient distribution of the queue length is known for the $M/M/\infty$ queue. Such a result is obtained by considering a two-dimensional Poisson point process representing the arrival time and the service time. A different approach is to bound the Entropy of the system (Chafaï, 2006). In this work, the author uses a diffusion approximation of the $M/M/\infty$ queuing system and uses the binomial-Poisson entropic inequalities to establish the rate of convergence.
@@
$M/M/n$ queue with $\alpha\in[1,\infty)$: In (Zeifman, 1991), the author uses the log-norm to characterize the convergence rate of the birth-and-death process, which applies to the $M/M/n$ setting. This method obtains the tight mixing rate of $(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ for $\alpha\geq 1$ and a sub-optimal mixing rate of $\mu-\frac{\lambda}{n-1}\ll(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ for $\alpha\in(1/2,1)$. The log-norm approach is also applied to analyze queueing models with queue-length-dependent admission control (Natvig, 1975), however, with a sub-optimal mixing rate (Van Doorn, 1981). This suggests that while simple, the log-norm approach is not strong enough to obtain sharp convergence rates for a wide variety of settings.
@@
$M/M/n$ queue with $\alpha=1/2$: The papers (Gamarnik and Goldberg, 2013; van Leeuwaarden and Knessl, 2011) uses the spectral representation of birth-and-death processes, previously established by (Karlin and McGregor, 1957), to obtain the rate of convergence of the $M/M/n$ queue in the Halfin-Whitt regime, i.e. $\alpha=1/2$. Most notably, the authors were able to characterize the phase transition of the mixing rate. In particular, as $\frac{n\mu-\lambda}{\sqrt{n}}$ increases, the mixing rate transitions from $(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ to $\approx 1$, establishing a phase shift from $M/M/1$-like behavior to $M/M/\infty$-like behavior. However, these results do not establish convergence in the Chi-square sense and are applicable only for $\alpha=1/2$.
$M/M/n/n+R$ queue: For a variant of $M/M/n$ with $n$ servers and $R$ waiting positions, (van Doorn and Zeifman, 2009; van Doorn, 2011) establish the rate of convergence to stationarity using the spectral representation method.
To the best of our knowledge, tight mixing results for $\alpha\in(0,1/2)$ and $\alpha\in(1/2,1)$ are still open prior to our work. And so, the overall mixing picture for $M/M/n$ is not complete. Our work fills the gap in the literature as we characterize tight mixing rates in the Super Halfin-Whitt regime and a tight mixing rate up to a constant for the Sub Halfin-Whitt regime. Moreover, (Gamarnik and Goldberg, 2013) is complementary to our work since this work focuses on the $\alpha=1/2$ regime while we obtain mixing results for every other regime, thus completing the mixing literature for $M/M/n$.
@@
Jackson networks: For Jackson networks, (Lorek and Szekli, 2015) shows exponential ergodicity but their bound using Cheeger’s inequality is not tight in terms of traffic, and the dependence of the mixing time on the spectral gap of the communication graph is not discussed. The spectral gap of the open Jackson network is established in (Mao and Xia, 2015) using the Markov chain decomposition method.
@@
Load-balancing systems: In addition to $M/M/n$ and Jackson networks, mixing results have been previously established for some load-balancing algorithms. The paper (Rutten and Mukherjee, 2023) obtained sub-geometric convergence for the spatial load-balancing system and (Ma and Maguluri, 2025) obtained $O(1/t)$ convergence rate for the join-the-shortest-queue system with $2$ servers. On the other hand, (Luczak and McDiarmid, 2006; Brightwell et al., 2018) obtained exponential convergence to a ball in the sub-Halfin-Whitt regime for the join-the-shortest-queue and power-of-$d$ policies.
1.3.1. Lyapunov drift method
@@
It is well-known that an irreducible Markov chain is positive recurrent if and only if it has a negative drift outside of a bounded set (Foster, 1953), which is famously known as the Foster-Lyapunov Theorem. Since then, several works have used the drift assumption to establish the convergence rate of the Markov chain. A long line of literature (Douc et al., 2018; Meyn and Tweedie, 2012, 1994; Down et al., 1995; Rosenthal, 1995; Lund et al., 1996; Baxendale, 2005; Taghvaei and Mehta, 2022; Hairer et al., 2011; Durmus and Moulines, 2014) establish geometric ergodicity for general Markov chains whenever we have the Lyapunov drift condition outside of some set and some control inside of that set (i.e. the set is a singleton or the set admits a minorization or a contraction condition). Recently, (Anderson et al., 2020) proposes a drift and hitting time approach instead of using the minorization assumption, and in a similar spirit, (Cattiaux et al., 2010) uses the exponential tail assumption of the hitting time to obtain Poincaré inequalities. However, these approaches do not allow us to obtain tight convergence rates for our case.
@@
The Lyapunov drift method is known to be very effective in establishing steady-state queueing results (Eryilmaz and Srikant, 2012; Hurtado-Lange and Maguluri, 2020; Jhunjhunwala et al., 2023) and beyond that, convergence and steady-state results in Stochastic Approximation and Optimization (Chen et al., 2021, 2019, 2020; Nguyen and Maguluri, 2023, 2024). As Lyapunov drift has been well-studied in the queuing literature, we naturally want to leverage this drift to prove mixing results. Previously, (Gurvich, 2014) obtained exponential convergence results for Markovian queues but the focus is not on obtaining tight mixing rates. For JSQ, (Braverman, 2020, 2023) uses the Lyapunov drift mixing framework by (Down et al., 1995). All of these works, however, only simply show exponential ergodicity and do not get a characterization on the mixing rate. This suggests that to get a tight mixing rate bound, a more fine-grained application of the Lyapunov drift method is required.
@@
In addition to the negative drift approach and the convergence in $\ell_{2}$ distance or TV distance, some other works have opted for convergence in Wasserstein space instead (Qu et al., 2023; Qin and Hobert, 2021; Durmus et al., 2015, 2023) since many systems such as the constant step size SGD may not converge in TV distance but does converge in Wasserstein distance (Qu et al., 2023). In this metric, (Qu et al., 2023) proposes the notion of contractive drift rather than the Foster-Lyapunov drift to obtain the finite-time convergence of various systems such as the single server queue and the constant step size SGD. Notably, this notion allows us to bypass the bounded set problem, where the drift and minorization approach is known to have limitations (Qin and Hobert, 2020). On the other hand, the approach of using a contraction condition to handle the bounded set (also known as drift and contraction) often leads to a tradeoff in terms of the drift rate and the bounded set size (Qu et al., 2023). Instead of using this approach, our work builds on the Lyapunov drift framework to obtain Chi-square convergence, and we argue that our work is another approach to handling the bounded set issue that is typically faced when using drift arguments to obtain mixing results.
@@
We summarize the literature and our results in Table 1. Our work covers the gap in the literature in the Light Traffic regime (where the arrival rate $\lambda_{n}$ is chosen such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$), the Sub-Halfin-Whitt regime (i.e. $\alpha\in(0,1/2)$), the Super-Halfin-Whitt regime (i.e. $\alpha\in(1/2,\infty)$) and the Heavy Traffic regime (i.e. $\alpha\in[1,\infty)$), completing the mixing literature for $M/M/n$. Additionally, note that our mixing results also match the asymptotic and the steady-state behavior of the system, that is the system behaves like $M/M/1$ for $\alpha\in(1/2,\infty)$ and $M/M/\infty$ for $\alpha\in(0,1/2)$.
1.4. Notations
@@
[…]
2. Model and Main Results
2.1. Model
@@ 2.2. Chi-square distance
(1) $\chi^{2}\left(P,Q\right)=\sum_{x\in\mathcal{S}}\frac{(P(x)-Q(x))^{2}}{Q(x)}.$
@@
(2) $\chi^{2}(P,Q)=\sum_{x\in\mathcal{S}}\frac{(P(x)-Q(x))^{2}}{Q(x)}=\sum_{x\in\mathcal{S}}Q(x)\left(\frac{P(x)^{2}}{Q(x)^{2}}-1\right)=\operatorname{Var}_{Q}\left(\frac{P}{Q}\right).$
@@
(3) $\chi^{2}(P,Q)=\sup_{g\in\mathcal{G}_{Q}}\frac{\left(\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]\right)^{2}}{\operatorname{Var}_{Q}[g(X)]}$
The proof of Proposition 1 is well-known and is presented in Appendix A.3. Additionally, Proposition 7.15 in (Polyanskiy and Wu, 2024) states that the Chi-square distance can be used to give an upper bound on the TV distance as well. From Proposition 1, we can substitute the function $g$ of our interests to obtain specific bounds. In particular, we obtain the following bound on the distance of the moments of the distributions by substituting $g(x)=x^{k}$.
@@
(4) $\left|\mathbb{E}_{P}\left[X^{k}\right]-\mathbb{E}_{Q}\left[X^{k}\right]\right|\leq\chi(P,Q)\sqrt{\mathbb{E}_{Q}\left[X^{2k}\right]-\mathbb{E}_{Q}\left[X^{k}\right]^{2}}.$
@@
(5) $\left|\mathbb{E}_{P}\left[e^{\theta X}\right]-\mathbb{E}_{Q}\left[e^{\theta X}\right]\right|\leq\chi\left(P,Q\right)\sqrt{\mathbb{E}_{Q}\left[e^{2\theta X}\right]-\mathbb{E}_{Q}\left[e^{\theta X}\right]^{2}}.$
The MGF bounds are particularly useful for obtaining exponential tail bounds. With the proper scaling, many classical queueing systems such as $M/M/n$ and JSQ possess an MGF (Hurtado-Lange and Maguluri, 2020; Jhunjhunwala et al., 2023), and thus one can derive exponential tail bounds for such systems.
@@
2.3. Main Results
Consider an $M/M/n$ queue with arrival rate $\lambda_{n}=n-n^{1-\alpha}$ for some $\alpha\in\mathbb{R}^{+}$ and service rate $\mu=1$ (unless stated otherwise). Such a parametrization of $\lambda$ is popularly known as the many-server-heavy-traffic regime and queueing systems are generally analyzed for $t,n\rightarrow\infty$. On the contrary, here we focus on a finite $t$ and a finite $n$. We obtain a bound on $\chi(\pi_{n,t},\nu_{n})$ of the form $e^{-\square t}$, thus, establishing the convergence of the finite-time queue length distribution $\pi_{n,t}$ to the stationary distribution $\nu_{n}$ as $t\rightarrow\infty$. Moreover, we provide a tight bound on the mixing rate, denoted by $\square$. This result is established in the following theorem.

### v4

@@ 1.3 Related works
Unlike many mixing time works which focus on the discrete-time finite state space Markov chains [34, 42, 6, 40, 85], analyzing the mixing behavior of queuing systems is a non-trivial endeavor as it involves handling an infinite, countable state space. Indeed, while there have been prior works on the mixing time of $M/M/n$ queues [67, 29, 86, 80] and other queuing systems [41, 49, 30, 7, 8, 43, 70], each of them can either only obtain tight finite-time bounds that match the asymptotic behavior only for a specific regime or setting, or can only show exponential ergodicity without an explicit rate [30, 7, 8], geometric convergence to a ball [43] or subgeometric convergence [70]. In particular, we shall provide a recap of mixing results for each of these regimes and settings.
@@
$M/M/1$ queue: The paper [54] is the first to characterize the transient behavior of an $M/M/1$ queue. Following this work, [73] uses the spectral representation of birth-and-death processes to obtain the rate of convergence. On the other hand, [67] uses a coupling argument to establish the distance to stationarity for the $M/M/1$ queue where the author upper bounds the distance to stationarity with the probability that the two processes will coalesce within time $T$, which can be further upper bounded by the probability of hitting 0 within time $T$. Unfortunately, both of these techniques are difficult to generalize to the $M/M/n$ queue given its more complicated dynamics.
@@
$M/M/\infty$ queue: The explicit characterization of the transient distribution of the queue length is known for the $M/M/\infty$ queue. Such a result is obtained by considering a two-dimensional Poisson point process representing the arrival time and the service time. A different approach is to bound the Entropy of the system [12]. In this work, the author uses a diffusion approximation of the $M/M/\infty$ queuing system and uses the binomial-Poisson entropic inequalities to establish the rate of convergence.
@@
$M/M/n$ queue: In [86], the author uses the log-norm to characterize the convergence rate of the birth-and-death process, which applies to the $M/M/n$ setting. This method obtains the tight mixing rate of $(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ for $\alpha\geq 1$ and a sub-optimal mixing rate of $\mu-\frac{\lambda}{n-1}\ll(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ for $\alpha\in(1/2,1)$. The log-norm approach is also applied to analyze queueing models with queue-length-dependent admission control [55], however, with a sub-optimal mixing rate [72]. This suggests that while simple, the log-norm approach is not strong enough to obtain sharp convergence rates for a wide variety of settings.
@@
In the Halfin-Whitt regime, [29] uses the spectral representation of birth-and-death processes, previously established by [36], to obtain the rate of convergence of the $M/M/n$ queue in the Halfin-Whitt regime, i.e. $\alpha=1/2$. Most notably, the authors were able to characterize the phase transition of the mixing rate. In particular, by letting the arrival rate $\lambda_{n}=n-B\sqrt{n}$ and $\mu=1$, the mixing rate transitions from $(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ to $\approx 1$ as $B$ increases from $0\rightarrow\infty$, establishing a phase transition from $M/M/1$-like behavior to $M/M/\infty$-like behavior even in the Halfin-Whitt regime. In their work, [29] establishes a bound on $|P_{ij}(t)-\pi_{j}(\infty)|$ for $|i-j|=O(\sqrt{n})$ and for a sufficiently large $n$, which only provides guarantee for sufficiently close states, but do not state finite-time moment bounds and tail bounds. These limitations motivate us to consider finite time bounds in the Chi-square metric in our work.
However, these prior results using the spectral representation can be used to obtain finite-time Chi-square bounds. While no prior work has explicitly stated such results, these results can be inferred by putting together several works in the literature. We obtain this and state a formal result in Appendix 12. We also present a discussion contrasting this work with our approach, as well as an in-depth comparison of our work with previous works on the $M/M/n$ systems. In particular, we would like to highlight that while our approach is applicable for any reversible Markov chain, the orthogonal polynomial approach is limited to birth-death chains.
@@
Jackson networks: For Jackson networks, [41] shows exponential ergodicity, but their bound using Cheeger’s inequality is not tight in terms of the traffic parameter, and the dependence of the mixing time on the spectral gap of the communication graph is not discussed. The spectral gap of the open Jackson network is established in [49] using the Markov chain decomposition method.
@@
Load-balancing systems: In addition to $M/M/n$ and Jackson networks, mixing results have been previously established for some load-balancing algorithms. The paper [70] obtained sub-geometric convergence for the spatial load-balancing system and [45] obtained $O(1/t)$ convergence rate for the join-the-shortest-queue system with $2$ servers. On the other hand, [43, 10] obtained exponential convergence to a ball in the sub-Halfin-Whitt regime for the join-the-shortest-queue and power-of-$d$ policies.
Birth and death processes: Besides queuing systems, birth and death processes (which entail $M/M/n$ queues) are common countable state space Markov chains. There are several works analyzing the spectral gap of birth and death processes [36, 73, 74, 72, 77, 75]. Additionally, it is known that these systems admit a system of orthogonal polynomials with respect to some spectral measure [19], which can be used to study exponential ergodicity [36, 29], $\ell_{2}$ convergence [14], or hitting time properties [76]. For non-homogeneous birth-death processes, [87] provides an analysis on the convergence of these processes and applies it to finite Markov chains. However, most of these works exploit the specific structure of birth and death processes (such as the spectral representation) which severely limits its applicability. In contrast, we aim to develop a more general framework to analyze the mixing behavior of reversible Markov chains.
1.3.1 Lyapunov drift method
@@
It is well-known that an irreducible Markov chain is positive recurrent if and only if it has a negative drift outside of a bounded set [28], which is famously known as the Foster-Lyapunov Theorem. Since then, several works have used the drift assumption to establish the convergence rate of the Markov chain. A long line of literature [22, 50, 51, 23, 68, 44, 5, 71, 31, 26] establish geometric ergodicity for general Markov chains whenever we have the Lyapunov drift condition outside of some set and some control inside of that set (i.e. the set is a singleton or the set admits a minorization or a contraction condition). Recently, [1] proposes a drift and hitting time approach instead of using the minorization assumption, and in a similar spirit, [11] uses the exponential integrability of the hitting time to obtain Poincaré inequalities. However, these approaches do not allow us to obtain tight convergence rates for our case.
@@
The Lyapunov drift method is known to be very effective in establishing steady-state queueing results [27, 33, 35] and beyond that, convergence and steady-state results in Stochastic Approximation and Optimization [16, 17, 15, 56, 57]. As Lyapunov drift has been well-studied in the queuing literature, we naturally want to leverage this drift to obtain mixing results. Previously, [30] obtained exponential convergence results for Markovian queues but the focus is not on obtaining tight mixing rates. For JSQ, [7, 8] uses the Lyapunov drift mixing framework by [23]. All of these works, however, only simply show exponential ergodicity and do not get a characterization on the mixing rate. This suggests that to get a tight mixing rate bound, a more fine-grained application of the Lyapunov drift method is required.
@@
In addition to the negative drift approach and the convergence in $\ell_{2}$ distance or TV distance, some other works have opted for convergence in Wasserstein space instead [65, 64, 25, 24] since many systems such as the constant step size SGD may not converge in TV distance but does converge in Wasserstein distance [65]. In this metric, [65] proposes the notion of contractive drift rather than the Foster-Lyapunov drift to obtain the finite-time convergence of various systems such as the single server queue and the constant step size SGD. Notably, this notion allows us to bypass the bounded set problem, where the drift and minorization approach is known to have limitations [63]. On the other hand, the approach of using a contraction condition to handle the bounded set (also known as drift and contraction) often leads to a tradeoff in terms of the drift rate and the bounded set size [65]. Instead of using this approach, our work builds on the Lyapunov drift framework to obtain Chi-square convergence, and we argue that our work is another approach to handling the bounded set issue that is typically faced when using drift arguments to obtain mixing results.
@@
We summarize our results and techniques in Table 1. Our work provides mixing time analysis in the Mean Field regime (where the arrival rate $\lambda_{n}$ is chosen such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$), the Sub-Halfin-Whitt regime (i.e. $\alpha\in(0,1/2)$) and the Super-Halfin-Whitt regime (i.e. $\alpha\in(1/2,\infty)$). Additionally, note that our mixing results also match the asymptotic and the steady-state behavior of the system, that is the system behaves like $M/M/1$ for $\alpha\in(1/2,\infty)$ and $M/M/\infty$ for $\alpha\in(0,1/2)$.
1.4 Notations
@@
In addition, for a well-behaved function $f$, we denote $\pi(f)=\mathbb{E}_{\pi}(f),\operatorname{Var}_{\pi}(f)=\mathbb{E}_{\pi}(f^{2})-\mathbb{E}_{\pi}(f)^{2}$, $\mathbf{1}$ as the vector of $1$s. For a set $K$, we denote $\pi(K)=\int_{K}d\pi$ and $1_{K}$ as the indicator function for the set $K$ where $1_{K}(x)=1$ if $x\in K$ and $1_{K}(x)=0$ otherwise. All logarithms in this paper are natural logarithms. Finally, we denote $\left[x\right]^{+}=\max\{0,x\}$ and we interpret $\frac{0}{0}$ as 0.
2 Model and Main Results
2.1 Model
@@ 2.2 Chi-square distance
$\chi^{2}\left(P,Q\right)=\sum_{x\in\mathcal{S}}\frac{(P(x)-Q(x))^{2}}{Q(x)}.$ (1)
@@
$\chi^{2}(P,Q)=\sum_{x\in\mathcal{S}}\frac{(P(x)-Q(x))^{2}}{Q(x)}=\sum_{x\in\mathcal{S}}Q(x)\left(\frac{P(x)^{2}}{Q(x)^{2}}-1\right)=\operatorname{Var}_{Q}\left(\frac{P}{Q}\right).$ (2)
@@
$\chi^{2}(P,Q)=\sup_{g\in\mathcal{G}_{Q}}\frac{\left(\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]\right)^{2}}{\operatorname{Var}_{Q}[g(X)]}$ (3)
The proof of Proposition 1 is well-known and is presented in Appendix 8.3. Additionally, Proposition 7.15 in [62] states that the Chi-square distance can be used to give an upper bound on the TV distance as well. From Proposition 1, we can substitute the function $g$ of our interests to obtain specific bounds. In particular, we obtain the following bound on the distance of the moments of the distributions by substituting $g(x)=x^{k}$.
@@
$\left|\mathbb{E}_{P}\left[X^{k}\right]-\mathbb{E}_{Q}\left[X^{k}\right]\right|\leq\chi(P,Q)\sqrt{\mathbb{E}_{Q}\left[X^{2k}\right]-\mathbb{E}_{Q}\left[X^{k}\right]^{2}}.$ (4)
@@
$\left|\mathbb{E}_{P}\left[e^{\theta X}\right]-\mathbb{E}_{Q}\left[e^{\theta X}\right]\right|\leq\chi\left(P,Q\right)\sqrt{\mathbb{E}_{Q}\left[e^{2\theta X}\right]-\mathbb{E}_{Q}\left[e^{\theta X}\right]^{2}}.$ (5)
The MGF bounds are particularly useful for obtaining exponential tail bounds. With the proper scaling, many classical queueing systems such as $M/M/n$ and JSQ possess an MGF [33, 35], and thus one can derive exponential tail bounds for such systems.
@@
2.3 Main Results
Consider an $M/M/n$ queue with arrival rate $\lambda_{n}=n-n^{1-\alpha}$ for some $\alpha\in\mathbb{R}^{+}$ and service rate $\mu=1$ (unless stated otherwise). Such a parametrization of $\lambda_{n}$ is popularly known as the many-server-heavy-traffic regime and queueing systems are generally analyzed for $t,n\rightarrow\infty$. On the contrary, here we focus on a finite $t\geq 0$ and a finite $n\geq 1$. We obtain a bound on $\chi(\pi_{n,t},\nu_{n})$ of the form $e^{-\square t}$, thus, establishing the convergence of the finite-time queue length distribution $\pi_{n,t}$ to the stationary distribution $\nu_{n}$ as $t\rightarrow\infty$. Moreover, we provide a tight bound on the mixing rate, denoted by $\square$. This result is established in the following theorem.